OpenStudy (anonymous):

MEDALLLLLLLLLSSSSSSSSSSSSSSSSSSSSSSSSSSSS!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Find the coordinates of the midpoint between point A(18, -6) and the origin. Answer Skip text editor options

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(18\quad ,&-6)\quad &(0\quad ,&0) \end{array}\\ \quad \\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\)

OpenStudy (anonymous):

wait do do I do next. I got 9,-6/2

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(18\quad ,&-6)\quad &(0\quad ,&0) \end{array}\\ \quad \\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\implies \left(\cfrac{18}{2}\quad ,\quad \cfrac{-6}{2} \right)\implies (9,-3) \) thus ( 9, -3 ) is the coordinates of the middle point of A and the origin (0,0)

OpenStudy (anonymous):

oh thanks you I didn't know what I did wrong, haha your the best

OpenStudy (anonymous):

how did you do that

OpenStudy (jdoe0001):

? what do you mean?

OpenStudy (anonymous):

nvm thanks you

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

wait can you help me with something else, it the same

OpenStudy (jdoe0001):

ok

OpenStudy (anonymous):

yea!!! ok. Find the coordinates of the midpoint of if A(-8, 7) and B(-9, 11).

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(-8\quad ,&7)\quad &(-9\quad ,&11) \end{array}\\ \quad \\ \text{middle point of 2 points }\\ \left(\cfrac{x_2 + x_1}{2}\quad ,\quad \cfrac{y_2 + y_1}{2} \right)\implies \left(\cfrac{-9+(-8)}{2}\quad ,\quad \cfrac{11+7}{2} \right)\)

OpenStudy (anonymous):

thanks that all I need

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

so its 8.5, 9 if I did it right

OpenStudy (jdoe0001):

-8.5, yes

OpenStudy (anonymous):

thanks again